# adF: a Novel System for Measuring Web Fingerprinting through Ads

### v2

@@ 9.3. Value frequency by attribute ($h$)
(3) $H(X)=-\sum_{i=1}^{n}P(x_{i})\cdot\log_{b}P(x_{i})$
Where $H$ is the entropy, $P$ is the probability of the appearance of the different values of the attribute, and $n$ is the total number of possible values for the attribute. The Shannon entropy is measured in bits and thus it uses a logarithm base equal to 2 ($b$ = 2).
To compare the entropy of attributes with different cardinality, we use a normalized version of Shannon’s entropy ($h$) dividing the equation 3 by ${\log_{2}(M)}$, where $M$ represents the number of values for each attribute (shannonentropy_sciencedirect, ).
(4) $h(X)=\frac{H(X)}{\log_{2}(M)}$
@@ 9.4. Summary of the analysis
Table 2 provides a summary of these three factors ($|N|$, $|S|$, $h$) for each attribute in the most relevant mobile app and browser-based configurations considered in this paper. Note that for attributes concatenated within a meta-attribute, we report the values for the meta-attribute.
The results in the table offer two main insights. On the one hand, we can assess that the configurations reported as most vulnerable to web fingerprinting in Section 8 are characterized by a higher number of informed attributes with high discrimination power (high cardinality and normalized entropy). For instance, the Chrome, Windows configuration, identified as highly vulnerable, reports 35 out of the 35 attributes listed in Table 2, with 8 of these attributes having a cardinality $\geq$ 299 and a normalized entropy $\geq$ 0,20. Contrary, the Firefox, Linux configuration reports 30 attributes, with just 1 exceeding the indicated cardinality and normalized entropy thresholds. Likewise, it is evident that mobile configurations generally report fewer attributes than desktop configurations. Furthermore, attributes’ cardinality and entropy tend to be lower in mobile configurations compared to desktop configurations.
On the other hand, Table 2 reveals those attributes with the most significant discrimination power (highlighted in bold). This represents very valuable information for browser and mobile apps developers, which can design solutions to mitigate the power discrimination of these attributes, reducing the vulnerability of the corresponding device configuration.
Figure 3. Performance of the proposed anti-fingerprinting solution represented by the reported $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ for each considered device configuration. The number over each bar represents the positive (blue) or negative (red) variation (in percentage) with respect to the same metric when ShieldF is not applied.
10. Countering web fingerprinting
The efficiency of fighting fingerprint is nowadays (and will be) an added value feature, which might positively affect the market share of a given browser. The browser industry is aware of this. As described in Section 2, they have (or are) developed(ing) different techniques to counter fingerprinting. These techniques, in essence, try to block the reporting of specific attributes or to reduce their cardinality and/or entropy. Our results in Section 8 indicate that certain browsers already implement quite efficient solutions (e.g., Safari or Firefox) while others (e.g., Chrome) provide an almost negligible defense against web fingerprinting.
@@ 10. Countering web fingerprinting
We propose ShieldF, a solution which relies on the most common technique already implemented by several browsers, this is, blocking/obfuscating the reporting of several attributes. This technique is simple, and browser developers are already using it, so it is likely that they are willing to adopt ShieldF. The key difference between ShieldF and existing solutions comes from the selection of the group of attributes to block. We propose to block those attributes which (1) have the largest discrimination power (i.e., highest cardinality and normalized entropy) and (2) do not affect (or have a very limited impact in) the end user experience [^4] For instance, an attribute revealing the screen size cannot be blocked even if its discrimination power is high. The screen size is required to adapt the content of a web-page to the screen size so it renders adequately.. To make an informed decision on the selected attributes to block, we rely on the results presented in Table 2, which reports the discrimination power (cardinality and normalized entropy) of all considered attributes in our study. In particular, we propose to block those attributes with a cardinality $>$ 25  [^5] Except for Bluetooth availability, since it is characterized by a binary value. and a normalized entropy $\geq$ 0,10, which we are certain do not affect the browsing experience. The 12 selected attributes are highlighted in bold in Table 2. Our proposal also applies to mobile apps with two exceptions: User Permissions state and Bluetooth availability attributes, because they are already unreported in mobile apps.
@@
10.1. Fingerprinting vulnerability reduction
@@
Figure 4. Comparison of the performance of ShieldF to the solutions proposed by Safari, Firefox, and Google. We report the $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ for the application of each of these solutions to a benchmark configuration, {Desktop, Windows, Chrome}.
10.2. Comparison to browsers’ anti-fingerprinting proposals
We compare the performance of ShieldF with solutions currently implemented by the main browsers. We consider the {Desktop, Chrome, Windows} device configuration as the benchmark. We apply the most protective version of the solution (e.g., activating opt-in anti-fingerprint options) implemented by Safari, Firefox and Chrome introduced in Section 2.3 as well as our own solution to this configuration.
@@ 10.2. Comparison to browsers’ anti-fingerprinting proposals
10.3. Comparison with countermeasure defenses
@@
While there is no definitive approach that can prevent web fingerprinting while maintaining the richness of a browser, the scientific community has developed various techniques and solutions aimed at mitigating the effects of web fingerprinting.
@@
Some works describe and analyze a specific method, such as Font fingerprinting (nikiforakis2015privaricator, ), Canvas fingerprinting (counterme_canvasDef, ; elbanna2019nonym, ; obidat2020canvas, ), or WebGL fingerprinting (wu2019rendered, ). Others take a broader view and quantify the combined use of functions and APIs most commonly employed for web fingerprinting (fiore2014countering, ; faizkhademi2015fpguard, ; laperdrix2015mitigating, ). While some solutions offer robust protection against browser fingerprinting, they often do so at the expense of usability (noscript, ; torproject_2022, ). The greatest challenge researchers face is providing complete coverage of modified attributes, as even minor discrepancies can make users more visible to trackers (boda2011user, ).
Figure 5 displays the values of $\mathcal{TV}$, $\mathcal{MV}$, and $\mathcal{A}$ for the configuration {Desktop, Chrome, Windows} as a benchmark, as well as our solution, ShieldF, for this configuration. The results indicate that in the general context of fingerprinting, solutions targeting a specific method like Canvas Defender (counterme_canvasDef, ), UniGL (wu2019rendered, ), or even FP-Random (laperdrix2017fprandom, ) —which focuses on defense techniques against Canvas fingerprinting, AudioContext fingerprinting, and unmasking browsers through JavaScript property ordering— offer protection that is seemingly marginal and makes almost no difference compared to not using them at all. Countermeasures like FP-Defense (ajay2022defense, ), FP-Block (torres2015fp, ), FP-Defender (moad2021fingerprint, ), and DCB (baumann2016disguised, ), which adopt a broader defense involving the combined use of JavaScript object-based functions and APIs, passive fingerprints, and HTTP headers, resulting in limited improvements in enhancing user privacy by preventing unwanted tracking. The FP-Defender solution would enhance resistance to web fingerprinting under the configuration of {Desktop, Chrome, Windows} by reducing $\mathcal{MV}$ ($\mathcal{TV}$) by approximately 11% (10%). However, this is still far from the improvement achieved by ShieldF, which is 60% (50%). On the other hand, the difference between $\mathcal{MV}$ and $\mathcal{TV}$ is larger with ShieldF compared to other solutions. This implies a lower accuracy of the adF fingerprinting system when our countermeasure ShieldF acts, as it has fewer attributes with high cardinality and entropy. This means that our adF fingerprinting system has a reduced capacity to identify unique fingerprints when ShieldF is active compared to other countermeasures.
One of the main challenges surrounding web fingerprinting and, thus, the development of defensive techniques is the continuous evolution of current web browser technologies. A proposed countermeasure can become obsolete in a matter of months, as each new browser version that modifies, adds, or even removes an object or API has a potential impact on the web fingerprinting ecosystem. For instance, the removal of access to NPAPI plugins brought the end of an important source of information for web fingerprinting and, consequently, for the mitigation efforts of some countermeasures (boda2011user, ; torres2015fp, ; faizkhademi2015fpguard, ; nikiforakis2015privaricator, ). In contrast, the later introduction of features such as Canvas API (counterme_canvasDef, ; elbanna2019nonym, ; obidat2020canvas, ), WebGL API (moad2021fingerprint, ; wu2019rendered, ), Battery Status API (ajay2022defense, ), and Web Audio API (laperdrix2017fprandom, ), revealed new vulnerabilities in web fingerprinting exploitation. As a result, the web fingerprints of the past are different from those we see today and will likely differ from those we will encounter in the coming years. Section 11 discusses the assessed countermeasures in more detail.
@@
Figure 5. Performance comparison of ShieldF with techniques and solutions aimed at mitigating the effects of web fingerprinting.. We report the $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ for the application of each of these solutions to a benchmark configuration, {Desktop, Windows, Chrome}. Dashed line represents the $\mathcal{MV}$ value of benchmark configuration {Desktop, Chrome, Windows}.
10.4. Empirical evaluation of ShieldF
@@
We implement ShieldF as an extension for any Chromium-based browser (Google Chrome, Microsoft Edge, Opera, Brave, etc). First of all, we prove that the 12 selected attributes highlighted in Table 2 are effectively blocked by the ShieldF extension. Moreover, 15 beta-testers have installed the ShieldF extension and used it during at least a month. They have been instructed to report any issue arising from their browsing experience. At the time of submitting this paper, none of the beta-testers have reported any major issue. We acknowledge that this initial evaluation cannot be considered exhaustive due to the limited number of beta-testers, but it offers initial empirical evidence that ShieldF has a negiglible impact in the browsing experience of users. The ShieldF extension is publicly available in the Chrome Web Store [^6] ShieldF is available on: https://chromewebstore.google.com/detail/shieldf/hdpnjiibilpldnnccggcpgilnpbdciia?hl=es&authuser=0 to allow any interested users to test it.
@@
-Summary: We conclude that a simple and easy-to-adopt solution by browsers and mobile app developers like ShieldF may substantially improve resilience against web fingerprinting. Note that our solution is based on a simple heuristic selection of attributes without the will of it being optimal, and thus, its performance can be considered a lower-bound improvement. For instance, each browser developer can take ShieldF code and the results reported in Table 2 as a starting point to define its optimized selection of attributes to be blocked in its browser.
@@
Table 3. Overview of large-scale fingerprinting studies: uniqueness (unq), no. collected attributes (No. att) and $H_{m}$.
Eckersley(eckersley2010, ) Laperdrix et al.(laperdrix2016beauty, ) Gómez-Boix et al.(gomezBoix2018hiding, ) Andriamilanto et al.(andriamilanto2021largescale, ) adF
@@
11. Related Work
11.1. Fingerprinting on the web
The profound privacy risks tied to web fingerprinting have motivated the research community to address the topic.
@@ 11.1. Fingerprinting on the web
The existing literature has extensively examined the feasibility of employing various JavaScript APIs for fingerprinting devices such as Canvas (mowery2012pixel, ; acar2014web, ; bursztein2016canvas, ; jiang2020tracking, ), WebGL API (cao2017cross, ; laor2022drawnapart, ), available fonts (nikiforakis2013cookieless, ; gomezBoix2018hiding, ), Web Audio API (englehardt2016online, ; cao2017cross, ; queiroz2019webaudio, ), Battery Status API (diaz2015leaking, ) or extensions (sjosten2017discovering, ; sanchez2017extension, ; starov2019unnecessarily, ; laperdrix2021fingerprinting, ; Konstantinos2022Chronos, ), as well as combinations of attributes; Mayer (mayer2009any, ) conducted, back in 2009, one of the pioneering works to uniquely identify individual devices. In particular, the author collected the content from navigator, screen, navigator.plugins, and navigator.mimeTypes attributes to compute the devices fingerprints. In addition, the research community has conducted different studies based on the measurement and web crawling techniques to assess to what extent web fingerprinting is used (nikiforakis2013cookieless, ; acar2013fpdetective, ; acar2014web, ; englehardt2016online, ; vastel2020fp, ; rizzo2021unveiling, ).
Most of the previous studies focus on desktop devices. However, the rapid growth of the mobile web and the development of its underlying technology has enabled the possibility to implement fingerprinting techniques to identify mobile devices uniquely (das2018web, ; kurtz2016fingerprinting, ; wu2016efficient, ; oliver2018fingerprintingthesis, ; heid2023functhooking, ; tiwarik2023androidhybridapps, ). Eubank et al. (eubank2013shining, ) conducted one of the earliest studies on mobile web tracking, analyzing cookies and HTTP headers used by browsers and revealing distinctions between the fingerprints of mobile and desktop devices. An overview of this literature can be found in the surveys by Laperdrix et al. (laperdrix2020fpSurvey, ) and Zhang et al. (zhang2022survey, )).
@@
In the context of our work, few existing papers have attempted to perform large-scale analysis to assess users’ exposure to web fingerprinting. A seminal work in the field by Eckersley (eckersley2010, ) performed the first large-scale analysis of web fingerprinting. By leveraging the Panopticlick (eckersley_panopticlick, ) website, the authors analyzed attributes collected from HTTP headers, JavaScript APIs, and Flash and Java plugins. More recently, Laperdrix et al. (laperdrix2016beauty, ) performed the first large-scale analysis considering fingerprints from both mobile and desktop devices. They based their study on data collected from the AmIUnique website (laperdrix_amiunique, ). These aforementioned works followed a similar approach, involving websites where volunteer users connect. These websites run scripts to compute the device’s fingerprint and assess the uniqueness of each fingerprint generated. Instead, Gómez-Boix et al. (gomezBoix2018hiding, ) and Andriamilanto et al. (andriamilanto2021largescale, ) collected fingerprints through a script deployed on one of the top 15 French websites, as per the Alexa traffic rank. Similar to (gomezBoix2018hiding, ; andriamilanto2021largescale, ), Li et al. (li2020wtmbf, ) deployed a fingerprinting tool on a real-world website and studied the effectiveness of web fingerprinting by collecting millions of samples.
@@
We present a novel approach for measuring web fingerprinting using ads as measurement venues rather than websites. This offers some advantages such as: 1) ability to define targeted measurement experiments for specific device configurations, which allows for first time to reveal the vulnerability at the granularity of individual device configurations; 2) independence from the willingness of volunteers or website owners to provide data; 3) capacity to audit the vulnerability to web fingerprinting in mobile apps for the first time.
Table 3 compares the major large-scale studies on web fingerprinting, including our own, across three dimensions: uniqueness, defined as the fraction of unique identified fingerprints for both desktop and mobile devices; number of collected attributes; and maximum diversity index. We observe that our methodology aligns well with the findings of the most recent study (andriamilanto2021largescale, ) in terms of uniqueness.
11.2. Fingerprinting countermeasures

### v3

@@ IX-C Value frequency by attribute ($h$)
$H(X)=-\sum_{i=1}^{n}P(x_{i})\cdot\log_{b}P(x_{i})$ (3)
Where $H$ is the entropy, $P$ is the probability distribution of the attribute’s values, and $n$ is the total possible values for the attribute. The Shannon entropy is measured in bits and thus it uses a logarithm base equal to 2 ($b$ = 2).
To compare the entropy of attributes with different cardinality, we use a normalized version of Shannon’s entropy ($h$) dividing the equation 3 by $\log_{2}(M)$}, where $M$ represents the number of values for each attribute.
$h(X)=\frac{H(X)}{\log_{2}(M)}$ (4)
@@ IX-D Summary of the analysis
Table II provides a summary of these three factors ($|N|$, $|S|$, $h$) for each attribute in the most relevant mobile app and browser-based configurations considered in this paper. Note that for attributes concatenated within a meta-attribute, we report the values for the meta-attribute.
The results in the table offer two main insights. On the one hand, we can assess that the configurations reported as most vulnerable to web fingerprinting in Section VII are characterized by a higher number of informed attributes with high discrimination power (high cardinality and normalized entropy). For instance, the Chrome, Windows configuration, identified as highly vulnerable, reports 35 out of the 35 attributes listed in Table II, with 8 of these attributes having a cardinality $\geq$ 299 and a normalized entropy $\geq$ 0.20. Contrary, the Firefox, Linux configuration reports 30 attributes, with just 1 exceeding the indicated cardinality and normalized entropy thresholds. Likewise, it is evident that mobile configurations generally report fewer attributes than desktop configurations. Furthermore, attributes’ cardinality and entropy tend to be lower in mobile configurations compared to desktop configurations.
On the other hand, Table II reveals those attributes with the most significant discrimination power (highlighted in bold). This represents very valuable information for browser and mobile apps developers, which can design solutions to mitigate the power discrimination of these attributes, reducing the vulnerability of the corresponding device configuration.
Remark. While this section focuses on static fingerprinting characteristics, Appendix C presents a longitudinal analysis of fingerprinting complexity over time, based on the joint entropy of full fingerprints across device configurations.
X Countering web fingerprinting
The efficiency of fighting fingerprint is nowadays (and will be) an added value feature, which might positively affect the market share of a given browser. The browser industry is aware of this. As described in Section II, they have (or are) developed(ing) different techniques to counter fingerprinting. These techniques, in essence, try to block the reporting of specific attributes or to reduce their cardinality and/or entropy. Our results in Section VII indicate that certain browsers already implement quite efficient solutions (e.g., Safari or Firefox) while others (e.g., Chrome) provide an almost negligible defense against web fingerprinting.
@@ X Countering web fingerprinting
Figure 3: The effectiveness of the proposed anti-fingerprinting solution is indicated by $\mathcal{MV}$, $\mathcal{TV}$ and $\mathcal{A}$ for each device configuration. Values above bars denote positive (blue) or negative (red) percentage variation compared to metrics without ShieldF.
We propose ShieldF, a solution which relies on the most common technique already implemented by several browsers, this is, blocking/obfuscating the reporting of several attributes. This technique is simple, and browser developers are already using it, so it is likely that they are willing to adopt ShieldF. The key difference between ShieldF and existing solutions comes from the selection of the group of attributes to block. We propose to block those attributes which (1) have the largest discrimination power (i.e., highest cardinality and normalized entropy) and (2) do not affect (or have a very limited impact in) the end user experience [^4] For instance, an attribute revealing the screen size cannot be blocked even if its discrimination power is high. The screen size is required to adapt the content of a web-page to the screen size so it renders adequately.. To make an informed decision on the selected attributes to block, we rely on the results presented in Table II, which reports the discrimination power (cardinality and normalized entropy) of all considered attributes in our study. In particular, we propose to block those attributes with a cardinality >> 25  [^5] Except for Bluetooth availability, since it is characterized by a binary value. and a normalized entropy $\geq$ 0.10, which we are certain do not affect the browsing experience. The 12 selected attributes are highlighted in bold in Table II. Note this selection also implicitly accounts for correlated signals: we only consider an attribute mitigated when it cannot be inferred through other unblocked sources. Our proposal also applies to mobile apps with two exceptions: User Permissions state and Bluetooth availability attributes, because they are already unreported in mobile apps.
@@
X-A Fingerprinting vulnerability reduction
@@
As a complementary check, we also quantify the information carried by the blocked attribute set via joint entropy. We compute the joint entropy of the same pooled dataset before and after applying ShieldF (formal definition in Appendix C). We observe a net drop of 2.15 bits (from 17.30 to 15.15), with an empirical $p$-value $10^{-4}$ from a permutation test under the null of random assignment.
X-B Comparison to browsers’ anti-fingerprinting proposals
We compare the performance of ShieldF with solutions currently implemented by the main browsers. We consider the {Desktop, Chrome, Windows} device configuration as the benchmark. We apply the most protective version of the solution (e.g., activating opt-in anti-fingerprint options) implemented by Safari, Firefox and Chrome introduced in Section II-C as well as our own solution to this configuration.
@@ X-B Comparison to browsers’ anti-fingerprinting proposals
X-C Comparison with countermeasure
@@
While there is no definitive approach that can prevent web fingerprinting while maintaining the richness of a browser, the scientific community has developed techniques and solutions aimed at mitigating the effects of web fingerprinting.
@@
Some works describe and analyze a specific method, such as Font fingerprinting [nikiforakis2015privaricator], Canvas fingerprinting [acar2014web], or WebGL fingerprinting [wu2019rendered]. Others take a broader view and quantify the combined use of functions and APIs most commonly employed for web fingerprinting [laperdrix2015mitigating]. While some solutions offer robust protection against browser fingerprinting, they often do so at the expense of usability [noscript]. The greatest challenge researchers face is providing complete coverage of modified attributes, as even minor discrepancies can make users more visible to trackers [boda2011user].
Figure 4: Comparing ShieldF to Safari, Firefox, and Google solutions. We assess $\mathcal{MV}$, $\mathcal{TV}$, and $\mathcal{A}$ on a benchmark setup, {Desktop, Windows, Chrome}.
@@
Figure 5: Comparing ShieldF with techniques to mitigate web fingerprinting effects. We assess $\mathcal{MV}$, $\mathcal{TV}$, and $\mathcal{A}$ on a benchmark setup, {Desktop, Windows, Chrome}. Dashed line denotes $\mathcal{MV}$ of benchmark configuration {Desktop, Chrome, Windows}.
Figure 5 displays the values of $\mathcal{TV}$, $\mathcal{MV}$, and $\mathcal{A}$ for the configuration {Desktop, Chrome, Windows} as a benchmark, as well as our solution, ShieldF, for this configuration. The results indicate that in the general context of fingerprinting, solutions targeting a specific method like Canvas Blocker [canvasbloquer_firefox], UniGL [wu2019rendered], or even FP-Random [laperdrix2017fprandom] —which focuses on defense techniques against Canvas fingerprinting, AudioContext fingerprinting, and unmasking browsers through JavaScript property ordering— offer protection that is seemingly marginal and makes almost no difference compared to not using them at all. Countermeasures like FP-Defense [ajay2022defense], FP-Block [torres2015fp], FP-Defender [moad2021fingerprint], and DCB [baumann2016disguised], which adopt a broader defense involving the combined use of JavaScript object-based functions and APIs, passive fingerprints, and HTTP headers, resulting in limited improvements in enhancing user privacy by preventing unwanted tracking. The FP-Defender solution would enhance resistance to web fingerprinting under the configuration of {Desktop, Chrome, Windows} by reducing $\mathcal{MV}$ ($\mathcal{TV}$) by approximately 11% (10%). However, this is still far from the improvement achieved by ShieldF, which is 60% (50%). On the other hand, the difference between $\mathcal{MV}$ and $\mathcal{TV}$ is larger with ShieldF compared to other solutions. This implies a lower accuracy of the adF fingerprinting system when our countermeasure ShieldF acts, as it has fewer attributes with high cardinality and entropy. This means that our adF fingerprinting system has a reduced capacity to identify unique fingerprints when ShieldF is active compared to other countermeasures.
One of the main challenges surrounding web fingerprinting and, thus, the development of defensive techniques is the continuous evolution of current web browser technologies. A proposed countermeasure can become obsolete in a matter of months, as each new browser version that modifies, adds, or even removes an object or API has a potential impact on the web fingerprinting ecosystem. For instance, the removal of access to NPAPI plugins brought the end of an important source of information for web fingerprinting and, consequently, for mitigation efforts of some countermeasures [boda2011user, torres2015fp, nikiforakis2015privaricator]. In contrast, the later introduction of features such as Canvas API [acar2014web], WebGL API [moad2021fingerprint, wu2019rendered], Battery Status API [ajay2022defense], and Web Audio API [laperdrix2017fprandom], revealed new vulnerabilities in web fingerprinting exploitation. As a result, the web fingerprints of the past are different from those we see today and will likely differ from those we will encounter in the coming years. Section 11 discusses assessed countermeasures in more detail.
X-D Empirical evaluation of ShieldF
@@
We implement ShieldF as an extension for any Chromium-based browser (Google Chrome, Microsoft Edge, Opera, Brave, etc). First of all, we prove that the 12 selected attributes highlighted in Table II are effectively blocked by the ShieldF extension. Moreover, 15 beta-testers have installed the ShieldF extension and used it for at least one month. They have been instructed to report any issue arising from their browsing experience. At the time of submitting this paper, none of the beta-testers have reported any major issue. The 15 beta-testers used desktop/laptop devices with up-to-date default installations and did not report using any additional privacy-preserving mechanisms beyond default browser settings. Their configurations were: {Chrome, Windows} 8/15 (53.3%), {Chrome, macOS} 5/15 (33.3%), and {Edge, Windows} 2/15 (13.3%). We acknowledge that this initial evaluation cannot be considered exhaustive due to the limited number of beta-testers, but it offers initial empirical evidence that ShieldF has a negligible impact on users’ browsing experience. The ShieldF extension is publicly available in the Chrome Web Store [^6] ShieldF is available at: https://chromewebstore.google.com/detail/shieldf/hdpnjiibilpldnnccggcpgilnpbdciia?hl=es&authuser=0 to allow any interested users to test it.
@@
-Summary: We conclude that a simple and easy-to-adopt solution by browsers and mobile app developers like ShieldF may substantially improve resilience against web fingerprinting. Note that our solution is based on a simple heuristic selection of attributes without the will of it being optimal, and thus, its performance can be considered a lower-bound improvement. Each browser developer can take ShieldF code and results reported in Table II as a starting point to define its optimized selection of attributes to be blocked in its browser.
@@
Table III: Overview of large-scale fingerprinting studies: uniqueness (unq), no. collected attributes (No. att) and $H_{m}$.
Eckersley[eckersley2010] Laperdrix et al.[laperdrix2016beauty] Gómez-Boix et al.[gomezBoix2018hiding] Andriamilanto et al.[andriamilanto2021largescale] adF
@@
XI Related Work
XI-A Fingerprinting on the web
@@
The profound privacy risks tied to web fingerprinting have motivated the research community to address the topic. The existing literature has extensively examined the feasibility of employing various JavaScript APIs for fingerprinting devices such as Canvas [acar2014web], WebGL API [cao2017cross], available fonts [gomezBoix2018hiding], Web Audio API [cao2017cross], Battery Status API [diaz2015leaking] or extensions [laperdrix2021fingerprinting], as well as combinations of attributes; Mayer [mayer2009any] conducted, back in 2009, one of the pioneering works to uniquely identify individual devices. In particular, the author collected the content from navigator, screen, navigator.plugins, and navigator.mimeTypes attributes to compute devices fingerprints. In addition, the research community has conducted different studies based on the measurement and web crawling techniques to assess to what extent web fingerprinting is used [acar2014web]. Most of the previous studies focus on desktop devices. However, the rapid growth of the mobile web and the development of its underlying technology has enabled the possibility to implement fingerprinting techniques to identify mobile devices uniquely [tiwarik2023androidhybridapps]. Eubank et al. [eubank2013shining] conducted one of the earliest studies on mobile web tracking, analyzing cookies and HTTP headers used by browsers and revealing distinctions between the fingerprints of mobile and desktop devices. A literature review can be found in the surveys by Laperdrix et al. [laperdrix2020fpSurvey] and Zhang et al. [zhang2022survey]).
@@
In the context of our work, few papers have attempted large-scale analysis to assess users’ exposure to web fingerprinting. A seminal work in the field by Eckersley [eckersley2010] performed the first large-scale analysis of web fingerprinting. By leveraging the Panopticlick [eckersley2010] website, the authors analyzed attributes collected from HTTP headers, JavaScript APIs, and Flash and Java plugins. More recently, Laperdrix et al. [laperdrix2016beauty] performed the first large-scale analysis considering fingerprints from both mobile and desktop devices. They based their study on data collected from the AmIUnique website [laperdrix2016beauty]. These aforementioned works followed a similar approach, involving websites where volunteer users connect. These websites run scripts to compute the device’s fingerprint and assess the uniqueness of each. Instead, Gómez-Boix et al. [gomezBoix2018hiding] and Andriamilanto et al. [andriamilanto2021largescale] collected fingerprints through a script deployed on one of the top 15 French websites, as per the Alexa traffic rank. Similar to [gomezBoix2018hiding, andriamilanto2021largescale], Li et al. [li2020wtmbf] deployed a fingerprinting tool on a real-world website and studied the effectiveness of web fingerprinting by collecting millions of samples.
@@
We present a novel approach for measuring web fingerprinting using ads as measurement venues rather than websites. This offers some advantages such as: 1) ability to define targeted measurement experiments for specific device configurations, which allows for first time to reveal the vulnerability at the granularity of individual device configurations; 2) independence from the willingness of volunteers or website owners to provide data; 3) capacity to audit the vulnerability to web fingerprinting in mobile apps for the first time. Table III compares the major large-scale studies on web fingerprinting, including our own, across three dimensions: uniqueness, defined as the fraction of unique identified fingerprints for both desktop and mobile devices; number of collected attributes; and maximum diversity index. We observe that our methodology aligns well with findings of the most recent study [andriamilanto2021largescale] in terms of uniqueness.
XI-B Fingerprinting countermeasures